# multimotion estimation problem MOT multiobject tracking

### v3

@@ 4.1.3 Pose-Velocity-Acceleration Estimator
The WNOA prior incentivizes locally-constant velocity, i.e., zero acceleration, which yields more physically plausible velocities, but is less accurate for motions with nonconstant velocity. This limitation is particularly important for objects that change direction, and a prior that can model nonzero accelerations is more representative of real-world motions. The WNOJ prior models the trajectory jerk as a continuous zero-mean, white noise Gaussian process (Tang et al., 2019),
@@
This is analogous to the pose-velocity estimator instead with a locally constant acceleration. The prior can model nonzero accelerations and is more applicable to real-world motions with objects that change direction. The pose-velocity-acceleration estimator state, $\mathbf{x}{}$, comprises the estimated pose transforms, $\mathscr{T}_{C}$, the body-centric velocities, $\left(\boldsymbol{\varpi}_{C_{k}}\right)_{k=1\dots K}$, the body-centric accelerations, $\left(\dot{\boldsymbol{\varpi}}_{C_{k}}\right)_{k=1\dots K}$. and the labeled landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{k}\coloneqq\left\{\mathbf{T}_{C_{k}C_{1}},\boldsymbol{\varpi}_{C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}}\right\}$ is defined as the transform, velocity, and acceleration at each time, $t_{k}$, and $\mathbf{x}{}_{jk}\coloneqq\left\{\mathbf{x}{}_{k},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of time steps and points in the label. The estimator proceeds analogously to the pose-velocity estimator using the same combined objective function in (21) with an expanded state transition function,
@@
The operating point is perturbed according to the transform perturbations, $\{\delta\boldsymbol{\xi}{k}\in\mathbb{R}^{6}\}$, velocity perturbations $\{\delta\boldsymbol{\varpi}_{k}\in\mathbb{R}^{6}\}$, acceleration perturbations $\{\delta\dot{\boldsymbol{\varpi}}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$, which are stacked to form the full state perturbation, $\delta\mathbf{x}{}$. The indicator matrices $\mathbf{P}_{jk}$ and $\mathbf{P}_{k}$ are defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ tuple and $\delta\mathbf{x}{}_{k}=\mathbf{P}_{k}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}}\right\}$ tuple. The optimal perturbation, $*\delta\mathbf{x}{}^{*}$, to minimize the linearized cost is the solution to the linear equation $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. The operating point is then updated using (14), (25), and
@@ 4.2 Geocentric Third-Party Estimation
Egomotion estimation techniques must be adapted to estimate the geocentric trajectories of dynamic objects that invalidate the static tracklet assumption. Initial versions of the ideas in Sections 4.2.1 and 4.2.2 were first published in Judd et al. (2018a) and Judd and Gammell (2020), respectively, and they are further developed here. The ideas in Section 4.2.3 are presented here for the first time. Pose-only estimates (Section 4.1.1) in an egocentric frame can be directly converted to a geocentric frame (Section 4.2.1). Inertial priors, such as the WNOA (Section 4.1.2) and WNOJ (Section 4.1.3) priors, are more complicated because bodies moving with a constant motion relative to a common inertial reference frame do not have constant motion relative to each other. Applying these motion priors to third-party trajectories requires rederiving the motion model in (6) and its associated Jacobians in (8) (Sections 4.2.2 and 4.2.3).
@@ 4.2.1 Pose-Only Estimator
$SE\left(3\right)$ transforms initially estimated using egomotion estimation techniques (Section 4.1.1) represent egomotion hypotheses calculated from the assumption that observed points are static. When the tracklets are moving, these estimates are the inverse of the objects’ motion relative to the sensor, i.e., the inverse of their egocentric motion. Egomotion hypothesis trajectories, ${{}^{\ell}{}}\mathscr{T}_{C}$, can be transformed into true third-party geocentric trajectories by compensating for the identified egomotion estimate, $\mathbf{T}_{C_{k}C_{1}}$, via
@@
where $\mathbf{F}_{{\ell_{k}}{\ell_{1}}}$ is the object deformation matrix and is assumed to be identity, (i.e., rigid-body motion). The initial sensor-to-object transform, $\mathbf{T}_{\ell{}_{1}C_{1}}$, relates the object frame from the sensor frame and is calculated blockwise from the initial translation and rotation. The initial translation, $\mathbf{p}^{C_{1}\ell{}_{1}}_{\ell{}_{1}}$, of the object frame is assigned to the observed centroid of the labeled points, $\mathscr{P}_{\ell{}}$, projected into the first observed frame,
@@
where $t_{j}$ is the first frame where the tracklet is observed. The initial rotation, $\mathbf{C}_{\ell_{1}C_{1}}$, is the initial frame orientation, which may be calculated from semantic information or be user specified (e.g., identity). Note that the egomotion estimate is only needed to convert third-party motions from the egocentric to geocentric frames and all motions are batch estimated first as egomotion hypotheses (Section 4.1.1) before identifying the egomotion label. This means motion-based heuristics and other application-specific information can be used to identify the egomotion before converting the remaining third-party motions to geocentric estimates.
@@ 4.2.2 Pose-Velocity Estimator
The WNOA prior is defined in an inertial frame, so third-party motions are estimated after identifying the egomotion. The techniques used to estimate the egomotion (Section 4.1.2) must be adapted to estimate third-party motions in a geocentric frame. The transform model, $\mathbf{z}$, in (6) is adjusted to transform egocentrically observed points through a geocentrically estimated state,
@@
The prior cost is the same as in (22). The linearized cost in (24) is then used to estimate the geocentric trajectory of every third-party motion in the scene. Note that the egomotion must be estimated before the third-party motions, so motion-based heuristics cannot be used to help identify the egomotion label as in pose-only estimation.
@@ 4.2.3 Pose-Velocity-Acceleration Estimator
The WNOJ prior is also defined in an inertial frame and is adapted using a similar procedure to the pose-velocity estimator (Section 4.2.2). The transform model, $\mathbf{z}$, in (6) takes the same form as (30). The motion model component of the measurement Jacobian, $\mathbf{G}_{jk}$, in (8) is given by the block row vector
@@
where $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ are defined in (31). The linearized cost takes the same form as (24) and uses the same prior cost term as (22). Note again that the egomotion must be estimated before the third-party motions and therefore motion-based heuristics cannot be used to identify the egomotion label.
@@ 5 Sliding-Window Multimotion Estimation
Sections 3 and 4 present a full-batch implementation of MVO. Full-batch estimation requires all observations of a given data segment at once but makes motion segmentation and estimation more accurate because motions are more distinct over longer periods of time. It can also reason about temporary occlusions after they occur, making it easier to determine when an object becomes occluded or unoccluded, as well as track motions through those occlusions. Full-batch estimation is computationally expensive and is often used in offline estimation systems. MVO can also be implemented as an online sliding-window pipeline that processes data as it is observed and is less computationally expensive than a full-batch implementation. Sliding-window estimation uses smaller batches of the $K$-most-recent observation frames, and maintaining consistent trajectories over consecutive batches presents several challenges. Motion segmentation is more difficult in a sliding-window pipeline because independent motions are often less distinct over short timescales. Motions must be associated across estimation windows to maintain label consistency because motions in separate windows are not guaranteed to be assigned the same label (Section 5.1). Unlike full-batch estimation, sliding-window estimation must reason about temporary occlusions as they occur. This requires motions to be extrapolated from, and interpolated between, direct observations (Section 5.2). Detecting when an object becomes unoccluded requires determining when a newly segmented motion is actually the reappearance of a previously occluded motion, i.e., motion closure (Section 5.3). Initial versions of these ideas were first presented in Judd and Gammell (2019b) and Judd and Gammell (2020), and they are further developed and fully detailed for the first time here.
@@ 5.1 Label Consistency
Sliding-window estimation requires consistently tracking motions across multiple windows. This involves associating motions in consecutive estimation windows to form a single trajectory over the entirety of the data. Each new window associates currently estimated motion labels, $\ell{}_{k}$, with previously estimated motion labels, $\ell{}_{k-1}^{\prime}$, based on the overlap in their labeled tracklets,
@@
Newly estimated motions without sufficient overlap with a previous motion are tracked as new object motions. Other motion- or appearance-based heuristics could also be employed to make this association more robust. The labeling and estimates from the current window are used to initialize the segmentation and egomotion hypotheses in the next window. Any current tracklets that also exist in the next window retain their labels, and the current trajectory estimates are extrapolated forward to the next time stamp to initialize the label set. This is done by extrapolating the geocentric estimates, $\mathbf{T}_{\ell{}_{k}\ell{}_{1}}$, forward and then converting them into egomotion hypotheses by rearranging (28),
@@ 5.2 Extrapolation and Interpolation
Unobserved motions can be estimated by modeling the expected motions of previous observations. This is less accurate than direct estimation, but can extend existing motions to initialize the next set of motion estimates as well as estimate motions through temporary occlusions. The accuracy of these estimates depends on the fidelity of the motion model to the true motion of the object, and inferring the motion of an object between two states, i.e., interpolation, is more accurate than from a single state, i.e., extrapolation. Pose extrapolation and interpolation require an estimate of the change in pose over time, i.e., the velocity. In the pose-only estimator (Sections 4.1.1 and 4.2.1), the velocity is discretely calculated from the relative transform between the final two poses,
@@
The body-centric velocity is continuously estimated in the pose-velocity (Sections 4.1.2 and 4.2.2) and pose-velocity-acceleration estimators (Sections 4.1.3 and 4.2.3). The pose-velocity-acceleration estimator also directly estimates the body-centric acceleration, which can be used for extrapolation and interpolation. This term is zero for the pose-only and pose-velocity estimators, i.e., constant-velocity extrapolation.
@@ 5.2.1 Extrapolation
where $\check{\cdot}$ denotes an extrapolated or interpolated value. Note that when the body-centric acceleration, $\dot{\boldsymbol{\varpi}}_{\ell{}_{k}}$, is zero, as with the pose-only and the pose-velocity estimators, (34) is mathematically equivalent to (16) and (17) with extra block zeros. The extrapolated local state is transformed to the global pose, velocity, and acceleration state via
@@
where the $\boldsymbol{\varpi}_{\check{\ell{}}}\left(\tau\right)$ term is omitted for the pose-only estimator and the $\dot{\boldsymbol{\varpi}}_{\check{\ell{}}}\left(\tau\right)$ term is omitted for both the pose-only and pose-velocity estimators. Estimates can be extrapolated forward or backward in time. The extrapolation accuracy will degrade over time, especially if the true motion exhibits significant changes in velocity or acceleration.
@@ 5.2.2 Interpolation
When the unobserved states are constrained by direct estimates on both sides, they can be estimated through interpolation. This is useful when tracking is resumed after an occlusion. The accuracy of the previously extrapolated estimates is improved because interpolation considers the directly observed data both before and after the occlusion. The pose-only state is linearly interpolated to an intermediate time, $t_{j}<\tau<t_{k}$, via
@@
which enforces a strict constant-velocity assumption. The pose-velocity and pose-velocity-acceleration states are interpolated (Anderson and Barfoot, 2015; Tang et al., 2019) via
@@
The block covariance matrix, $\mathbf{Q}_{j}\left(\tau\right)$, is given by (23) for the WNOA prior and (27) for the WNOJ prior. The interpolated state is then transformed to the global state via (35). Calculating the interpolation involves estimates both before and after the occlusion. This requires recognizing when a newly discovered motion can be explained by the reappearance of a previously occluded object, i.e., motion closure (Section 5.3).

### v4

@@ 4.1.3 Pose-Velocity-Acceleration Estimator
The WNOA prior incentivizes locally-constant velocity, i.e., zero acceleration, which yields more physically plausible velocities, but is less accurate for motions with nonconstant velocity. This limitation is particularly important for objects that change direction, and a prior that can model nonzero accelerations is more representative of real-world motions.
The WNOJ prior models the trajectory jerk as a continuous zero-mean, white noise Gaussian process (Tang et al., 2019),
@@
This is analogous to the pose-velocity estimator instead with a locally constant acceleration. The prior can model nonzero accelerations and is more applicable to real-world motions with objects that change direction.
The pose-velocity-acceleration estimator state, $\mathbf{x}{}$, comprises the estimated pose transforms, $\mathscr{T}_{C}$, the body-centric velocities, $\left(\boldsymbol{\varpi}_{C_{k}}\right)_{k=1\dots K}$, the body-centric accelerations, $\left(\dot{\boldsymbol{\varpi}}_{C_{k}}\right)_{k=1\dots K}$. and the labeled landmark points, $\left\{\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}_{j=1\dots|\mathscr{P}_{C}|}$. The state $\mathbf{x}{}_{k}\coloneqq\left\{\mathbf{T}_{C_{k}C_{1}},\boldsymbol{\varpi}_{C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}}\right\}$ is defined as the transform, velocity, and acceleration at each time, $t_{k}$, and $\mathbf{x}{}_{jk}\coloneqq\left\{\mathbf{x}{}_{k},\mathbf{p}^{{j_{1}}{C_{1}}}_{C_{1}}\right\}$ is defined for each pair of time steps and points in the label.
The estimator proceeds analogously to the pose-velocity estimator using the same combined objective function in (13) with an expanded state transition function,
@@
The operating point is perturbed according to the transform perturbations, $\{\delta\boldsymbol{\xi}{k}\in\mathbb{R}^{6}\}$, velocity perturbations $\{\delta\boldsymbol{\varpi}_{k}\in\mathbb{R}^{6}\}$, acceleration perturbations $\{\delta\dot{\boldsymbol{\varpi}}_{k}\in\mathbb{R}^{6}\}$, and landmark perturbations, $\{\delta\boldsymbol{\zeta}_{j}\in\mathbb{R}^{3}\}$, which are stacked to form the full state perturbation, $\delta\mathbf{x}{}$. The indicator matrices $\mathbf{P}_{jk}$ and $\mathbf{P}_{k}$ are defined such that $\delta\mathbf{x}{}_{jk}=\mathbf{P}_{jk}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}},\mathbf{p}^{j_{1}C_{1}}_{\mathrm{op},C_{1}}\right\}$ tuple and $\delta\mathbf{x}{}_{k}=\mathbf{P}_{k}\delta\mathbf{x}{}$ is the perturbation of the $\left\{\mathbf{T}_{\mathrm{op},C_{k}C_{1}},\boldsymbol{\varpi}_{\mathrm{op},C_{k}},\dot{\boldsymbol{\varpi}}_{C_{k}}\right\}$ tuple.
The optimal perturbation, $\delta\mathbf{x}{}^{*}$, to minimize the linearized cost is the solution to the linear equation $\mathbf{A}\delta\mathbf{x}{}^{*}=\mathbf{b}$. The operating point is then updated using (9), (17), and
@@ 4.2 Geocentric Third-Party Estimation
Egomotion estimation techniques must be adapted to estimate the geocentric trajectories of dynamic objects that invalidate the static tracklet assumption. Initial versions of the ideas in Sections 4.2.1 and 4.2.2 were first published in Judd et al. (2018a) and Judd and Gammell (2020), respectively, and they are further developed here. The ideas in Section 4.2.3 are presented here for the first time.
Pose-only estimates (Section 4.1.1) in an egocentric frame can be directly converted to a geocentric frame (Section 4.2.1). Inertial priors, such as the WNOA (Section 4.1.2) and WNOJ (Section 4.1.3) priors, are more complicated because bodies moving with a constant motion relative to a common inertial reference frame do not have constant motion relative to each other. Applying these motion priors to third-party trajectories requires rederiving the motion model in (6) and its associated Jacobians in (8) (Sections 4.2.2 and 4.2.3).
@@ 4.2.1 Pose-Only Estimator
$SE\left(3\right)$ transforms initially estimated using egomotion estimation techniques (Section 4.1.1) represent egomotion hypotheses calculated from the assumption that observed points are static. When the tracklets are moving, these estimates are the inverse of the objects’ motion relative to the sensor, i.e., the inverse of their egocentric motion.
Egomotion hypothesis trajectories, ${{}^{\ell}{}}\mathscr{T}_{C}$, can be transformed into true third-party geocentric trajectories by compensating for the identified egomotion estimate, $\mathbf{T}_{C_{k}C_{1}}$, via
@@
where $\mathbf{F}_{{\ell_{k}}{\ell_{1}}}$ is the object deformation matrix and is assumed to be identity, (i.e., rigid-body motion).
The initial sensor-to-object transform, $\mathbf{T}_{\ell{}_{1}C_{1}}$, relates the object frame from the sensor frame and is calculated blockwise from the initial translation and rotation. The initial translation, $\mathbf{p}^{C_{1}\ell{}_{1}}_{\ell{}_{1}}$, of the object frame is assigned to the observed centroid of the labeled points, $\mathscr{P}_{\ell{}}$, projected into the first observed frame,
@@
where $t_{j}$ is the first frame where the tracklet is observed. The initial rotation, $\mathbf{C}_{\ell_{1}C_{1}}$, is the initial frame orientation, which may be calculated from semantic information or be user specified (e.g., identity).
Note that the egomotion estimate is only needed to convert third-party motions from the egocentric to geocentric frames and all motions are batch estimated first as egomotion hypotheses (Section 4.1.1) before identifying the egomotion label. This means motion-based heuristics and other application-specific information can be used to identify the egomotion before converting the remaining third-party motions to geocentric estimates.
@@ 4.2.2 Pose-Velocity Estimator
The WNOA prior is defined in an inertial frame, so third-party motions are estimated after identifying the egomotion. The techniques used to estimate the egomotion (Section 4.1.2) must be adapted to estimate third-party motions in a geocentric frame.
The transform model, $\mathbf{z}$, in (6) is adjusted to transform egocentrically observed points through a geocentrically estimated state,
@@
The prior cost is the same as in (14). The linearized cost in (16) is then used to estimate the geocentric trajectory of every third-party motion in the scene.
Note that the egomotion must be estimated before the third-party motions, so motion-based heuristics cannot be used to help identify the egomotion label as in pose-only estimation.
@@ 4.2.3 Pose-Velocity-Acceleration Estimator
The WNOJ prior is also defined in an inertial frame and is adapted using a similar procedure to the pose-velocity estimator (Section 4.2.2).
The transform model, $\mathbf{z}$, in (6) takes the same form as (22). The motion model component of the measurement Jacobian, $\mathbf{G}_{jk}$, in (8) is given by the block row vector
@@
where $\mathbf{Z}_{1}$ and $\mathbf{Z}_{3}$ are defined in (23). The linearized cost takes the same form as (16) and uses the same prior cost term as (14).
Note again that the egomotion must be estimated before the third-party motions and therefore motion-based heuristics cannot be used to identify the egomotion label.
@@ 5 Sliding-Window Multimotion Estimation
Sections 3 and 4 present a full-batch implementation of MVO. Full-batch estimation requires all observations of a given data segment at once but makes motion segmentation and estimation more accurate because motions are more distinct over longer periods of time. It can also reason about temporary occlusions after they occur, making it easier to determine when an object becomes occluded or unoccluded, as well as track motions through those occlusions. Full-batch estimation is computationally expensive and is often used in offline estimation systems.
MVO can also be implemented as an online sliding-window pipeline that processes data as it is observed and is less computationally expensive than a full-batch implementation. Sliding-window estimation uses smaller batches of the $K$-most-recent observation frames, and maintaining consistent trajectories over consecutive batches presents several challenges.
Motion segmentation is more difficult in a sliding-window pipeline because independent motions are often less distinct over short timescales. Motions must be associated across estimation windows to maintain label consistency because motions in separate windows are not guaranteed to be assigned the same label (Section 5.1). Unlike full-batch estimation, sliding-window estimation must reason about temporary occlusions as they occur. This requires motions to be extrapolated from, and interpolated between, direct observations (Section 5.2). Detecting when an object becomes unoccluded requires determining when a newly segmented motion is actually the reappearance of a previously occluded motion, i.e., motion closure (Section 5.3). Initial versions of these ideas were first presented in Judd and Gammell (2019b) and Judd and Gammell (2020), and they are further developed and fully detailed for the first time here.
@@ 5.1 Label Consistency
Sliding-window estimation requires consistently tracking motions across multiple windows. This involves associating motions in consecutive estimation windows to form a single trajectory over the entirety of the data.
Each new window associates currently estimated motion labels, $\ell{}_{k}$, with previously estimated motion labels, $\ell{}_{k-1}^{\prime}$, based on the overlap in their labeled tracklets,
@@
Newly estimated motions without sufficient overlap with a previous motion are tracked as new object motions. Other motion- or appearance-based heuristics could also be employed to make this association more robust.
The labeling and estimates from the current window are used to initialize the segmentation and egomotion hypotheses in the next window. Any current tracklets that also exist in the next window retain their labels, and the current trajectory estimates are extrapolated forward to the next time stamp to initialize the label set. This is done by extrapolating the geocentric estimates, $\mathbf{T}_{\ell{}_{k}\ell{}_{1}}$, forward and then converting them into egomotion hypotheses by rearranging (20),
@@ 5.2 Extrapolation and Interpolation
Unobserved motions can be estimated by modeling the expected motions of previous observations. This is less accurate than direct estimation, but can extend existing motions to initialize the next set of motion estimates as well as estimate motions through temporary occlusions. The accuracy of these estimates depends on the fidelity of the motion model to the true motion of the object, and inferring the motion of an object between two states, i.e., interpolation, is more accurate than from a single state, i.e., extrapolation.
Pose extrapolation and interpolation require an estimate of the change in pose over time, i.e., the velocity. In the pose-only estimator (Sections 4.1.1 and 4.2.1), the velocity is discretely calculated from the relative transform between the final two poses,
@@
The body-centric velocity is continuously estimated in the pose-velocity (Sections 4.1.2 and 4.2.2) and pose-velocity-acceleration estimators (Sections 4.1.3 and 4.2.3).
The pose-velocity-acceleration estimator also directly estimates the body-centric acceleration, which can be used for extrapolation and interpolation. This term is zero for the pose-only and pose-velocity estimators, i.e., constant-velocity extrapolation.
@@ 5.2.1 Extrapolation
where $\check{\cdot}$ denotes an extrapolated or interpolated value. Note that when the body-centric acceleration, $\dot{\boldsymbol{\varpi}}_{\ell{}_{k}}$, is zero, as with the pose-only and the pose-velocity estimators, (26) is mathematically equivalent to (11) and (12) with extra block zeros.
The extrapolated local state is transformed to the global pose, velocity, and acceleration state via
@@
where the $\boldsymbol{\varpi}_{\check{\ell{}}}\left(\tau\right)$ term is omitted for the pose-only estimator and the $\dot{\boldsymbol{\varpi}}_{\check{\ell{}}}\left(\tau\right)$ term is omitted for both the pose-only and pose-velocity estimators.
Estimates can be extrapolated forward or backward in time. The extrapolation accuracy will degrade over time, especially if the true motion exhibits significant changes in velocity or acceleration.
@@ 5.2.2 Interpolation
When the unobserved states are constrained by direct estimates on both sides, they can be estimated through interpolation. This is useful when tracking is resumed after an occlusion. The accuracy of the previously extrapolated estimates is improved because interpolation considers the directly observed data both before and after the occlusion.
The pose-only state is linearly interpolated to an intermediate time, $t_{j}<\tau<t_{k}$, via
@@
which enforces a strict constant-velocity assumption.
The pose-velocity and pose-velocity-acceleration states are interpolated (Anderson and Barfoot, 2015; Tang et al., 2019) via
@@
The block covariance matrix, $\mathbf{Q}_{j}\left(\tau\right)$, is given by (15) for the WNOA prior and (19) for the WNOJ prior. The interpolated state is then transformed to the global state via (27).
Calculating the interpolation involves estimates both before and after the occlusion. This requires recognizing when a newly discovered motion can be explained by the reappearance of a previously occluded object, i.e., motion closure (Section 5.3).